A 10 kg mass with an initial velocity of 6 i m/s collides and sticks to a 5 kg mass with an initial velocity of - 5 j m/s. Find the direction (in degrees) of the final velocity of the two masses with respect to the direction of the x axis. (Note: take your calculated angle as positive.)
Answer #1

here,

mass of object 1 , m1 = 10 kg

initial velocity of 1 , u1 = 6 i m/s

mass of object 2 , m2 = 5 kg

initial velocity of 2 , u2 = - 5 j m/s

let the final velocity be v

using conservation of momentum

m1 * u1 + m2 * u2 = (m1 + m2) * v

10 * 6 i - 5 * 5 j = ( 10 + 5) * v

solving for v

v = 4 i m/s - 1.67 j m/s

the direction of final velocity with respect to direction of x-axis , theta = arctan((-1.67)/4)

theta = 22.7 degree